Lehrstuhl A für Mathematik Aachen, den 20.12.2001

Prof. Dr. R. L. Stens, Tel.: 80-94532 Dipl.-Math. B. Ohligs, Tel.: 80-94315

9. Übung zu Zahlbereichserweiterungen

(Abgabe: Donnerstag, 10.01.2002, \underline{vor} der Übung oder bis 10 Uhr im Übungskasten vor dem Sekretariat des Lehrstuhls)

Aufgabe 1: Seien a ∈ ℤ, n ∈ ℕ, n ≥ 2. Zeigen Sie:

a) Gilt x^{n} = a für ein x ∈ ℚ, so ist x ∈ ℤ.  3

b) Die Gleichung x^{n} = a hat genau dann eine Lösung in ℚ, wenn a = m^{n} für ein

m ∈ ℤ.  1

Aufgabe 2: Zeigen sie, dass es keinen Isomorphismus zwischen der additiven Gruppe

(ℚ, +) und der multiplikativen Gruppe (ℚ^{+}, ·) gibt.  4

Aufgabe 3: Bestimmen Sie alle Ringhomomorphismen φ : ℚ → ℝ und ψ : ℝ → ℚ.  5

Aufgabe 4: Es sei G die Menge aller Abbildungen f : ℤ → ℤ, zu denen es eine natürliche Zahl c gibt, so dass für alle x, y ∈ ℤ gilt

|f(x + y) − f(x) − f(y)| ≤ c.

Für f, g ∈ G sei f ⊕ g : ℤ → ℤ definiert durch (f ⊕ g)(x) := f(x) + g(x) für alle x, y ∈ ℤ. Zeigen Sie:

a) (G, ⊕) ist eine kommutative Gruppe.  2
b) |f(mn) − nf(m)| ≤ (n − 1)c für alle f ∈ G und m, n ∈ ℕ.  3

c) Für jedes f ∈ G existiert

Φ(f) := lim_{n→∞} \frac{f(n)}{n} ∈ ℝ. 3

d) Die in c) angegebene Abbildung liefert einen surjektiven Gruppenhomomorphismus Φ : (G, ⊕) → (ℝ, +).

Hinweis: Für r ∈ ℝ betrachte man die Abbildung ℤ → ℤ, x ↦ [rx].  4

Aufgabe 5: Eine Folge (a_{n})_{n} sei definiert durch a_{1} := α mit 0 ≤ α ≤ \frac{1}{2} und a_{n+1} := a_{n}^{2} + \frac{1}{4} für alle n ∈ ℕ. Zeigen Sie:

a) (a_{n})_{n} ist monoton wachsend.  1
b) Für alle n ∈ ℕ ist 0 ≤ a_{n} ≤ \frac{1}{2}.  1
c) (a_{n})_{n} konvergiert.  1
d) Bestimmen Sie lim_{n→∞} a_{n}.  2

30

Augustin Louis Cauchy

(Teil 2) In 1815 Cauchy lost out to Binet for a mechanics chair at the Ecole Polytechnique, but then was appointed assistant professor of analysis there. He was responsible for the second year course. In 1816 he won the Grand Prix of the French Academy of Science for a work on waves. He achieved real fame however when he submitted a paper to the Institute solving one of Fermat’s claims on polygonal numbers made to Mersenne. Politics now helped Cauchy into the Academy of Sciences when Carnot and Monge fell from political favour and were dismissed and Cauchy filled one of the two places.

In 1817 when Biot left Paris for an expedition to the Shetland Islands in Scotland Cauchy filled his post at the Collège de France. There he lectured on methods of integration which he had discovered, but not published, earlier. Cauchy was the first to make a rigorous study of the conditions for convergence of infinite series in addition to his rigorous definition of an integral. His text Cours d’analyse in 1821 was designed for students at Ecole Polytechnique and was concerned with developing the basic theorems of the calculus as rigorously as possible. He began a study of the calculus of residues in 1826 in Sur un nouveau genre de calcul analogue au calcul infinétesimal while in 1829 in Leçons sur le Calcul Différential he defined for the first time a complex function of a complex variable.

Cauchy did not have particularly good relations with other scientists. His staunchly Catholic views had him involved on the side of the Jesuits against the Académie des Sciences. He would bring religion into his scientific work as for example he did on giving a report on the theory of light in 1824 when he attacked the author for his view that Newton had not believed that people had souls. (Fortsetzung folgt)